# Teaching Spell Checkers to Teach: Pedagogical Program Synthesis for Interactive Learning

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
Figure 1. This figures illustrates how in SPIRE identifies misspellings (left), generates a program for intervention (center) and subsequently renders them in the interface (right) for learner’s inquiry to surface word related insights
@@ 1. Introduction
School-age students often learn spelling by memorization (hegland2021beneath). However, the way in which a word is spelled is more than just a convention of letters. It reflects patterns of meaning, sound, structure, and spelling conventions that connect words to one another and to their histories. For instance, consider the word ‘reach’ meaning to extend or stretch out. A majority of us can spell this word with ease, and are unlikely to spell it as say <reech>, in which the long /i/ sound is mapped to the grapheme <ee>. In fact, English represents the /i/ sound in many ways, including <ee> as in see, <ea> as in reach, <ie> as in piece, and <e-e> as in these. Memorization helps us recall the correct form. However, for many learners, especially children with language-based learning disorders, rote memorization is insufficient because it recruits only one aspect of the lexical quality of the word (phonology or orthography/spelling) (perfetti2007reading; seidenberg1989distributed). What supports their learning is understanding the logic behind these alternatives; for example, noticing that ‘reach’ and ‘teach’ both use the <ea> to represent /iː/, and that this pattern traces back to the Old English reccan (etymonline_reach). Seen this way, spelling becomes less an arbitrary code to memorize and more a system of patterns that can be reasoned about, making the language more learnable and meaningful (tbox2). In fact, learners who struggle with spelling through memorization are often referred to speech and language therapy. In therapy, speech-language pathologists (SLPs) use inquiry as an approach to help learners make sense of how words are spelled and built. For example, if a learner writes <reech> for “reach”, the SLP may begin by asking how the learner decided on that spelling. If the learner explains that /i/ is often spelled <ee> as in “see” or “speech”, the SLP can validate this reasoning, then introduce a contrast, such as “teach” or “teacher”, to highlight the alternative <ea> spelling. With a different error, such as using <sine> for “sign”, the SLP may draw attention to the silent <g> and connect it to morphologically related words like “signal” or “signature”. This practice, known as Structured Word Inquiry (SWI) (effects_bowers_2010), turns spelling into a process of investigation rather than memorization, drawing on phonology (the relationship between sounds and letters), semantics (meaning of the word), morphology (the structure and meaning of word parts such as roots, prefixes, and suffixes), and etymology (the historical origins and evolution of words) to explain why words are spelled the way they are. Through these inquiry-based interventions, SLPs help learners build durable knowledge that extends beyond a single correction to a broader understanding of how written language works. Consequently, they learn how to independently investigate words.
@@
In contrast, everyday writing tools take a much narrower approach to spelling errors. A misspelling is detected, underlined, and replaced with the most probable correction. Whether a learner types “reech” because they overgeneralized a phonics rule, misapplied a morphological pattern, or simply hit the wrong key, the system’s response is the same. This transactional design supports error-free text, but it bypasses the chance to engage learners in noticing patterns or testing hypotheses. This is a missed opportunity because therapy time is scarce (exploring_lewis_2025) and spelling occurs within these tools in the context of authentic writing tasks, such as science reports and essay writing. Without support in these everyday contexts, the carryover of SWI pedagogy from therapy to classroom writing is limited. What is needed is a way to embed the principles of SWI into the tools learners already use, turning each correction into a brief, situated learning experience. This motivates our central research question: How might we bring the rich, inquiry-based pedagogy of SLPs into everyday writing tools such that spelling errors become an opportunity to learn about the writing system?
@@
Our key contributions include: (1) a Pedagogical Program Synthesis that represents SLP’s instructional moves in a domain-specific language and assembles short, conditional programs from learner errors in real time, (2) SPIRE, a spell-checker that embeds these synthesized programs into lightweight, interactive interfaces, and (3) an empirical evaluation with experts and learners demonstrating how SPIRE aligns with professional practice and integrates into writing workflows.
@@ 2. Related Work
Our work builds on three interconnected bodies of research: research on spelling development and learning difficulties in school-age students, evidence-based interventions designed to support struggling spellers, and recent efforts leveraging AI to deliver adaptive literacy instruction.
@@
Vellutino et al. (specific_vellutino_2004) synthesized four decades of empirical work, establishing that difficulties in phonological awareness, rapid naming, and verbal working memory constitute core deficits that persist in development. Longitudinal studies of upper elementary students with reading or spelling disabilities reveal how these difficulties in phonological information processing evolve throughout formal schooling (developmental_schmidt_2021). However, the presentation of symptoms varies considerably between individuals, requiring adaptive intervention approaches (developmental_kuerten_2019). Corpus-based documentation allowed for an empirical understanding of characteristic dyslexic error types. DysList systematically cataloged error patterns in multiple languages (dyslist_rello_2014), while subsequent analysis identified specific phoneme-grapheme correspondence failures that persist despite repeated exposure to high-frequency vocabulary (analysis_cidrim_2024). These difficulties extend beyond traditional literacy contexts into digital environments, where searchers with dyslexia exhibit distinct information-seeking behaviors requiring specialized interface considerations (understanding_morris_2018).
@@
Ehri’s phase model (phases_ehri_1995; learning_ehri_2005) describes progression through distinct developmental stages: pre-alphabetic, partial alphabetic, full alphabetic, and consolidated alphabetic, each characterized by qualitatively different approaches to word recognition. This framework establishes grapheme-phoneme knowledge as the basis for reading acquisition in alphabetic languages (graphemephoneme_ehri_1998). However, Silliman et al. (spelling_silliman_2006) demonstrated that preadolescents with atypical language skills show distinct error patterns reflecting deficits in phonological, morphological, and orthographic dimensions, indicating that effective intervention must address multiple representational systems simultaneously.
@@
Müller et al. (syllablebased_mller_2020) found that syllable-based reading instruction produces significant improvements in both word-reading accuracy and comprehension in second-grade students.
@@
Meta-analytical findings demonstrate more severe reading impairments in deep orthographies such as English compared to shallow orthographies with consistent grapheme-phoneme mappings (orthographic_carioti_2021), suggesting that intervention approaches must account for the characteristics of the writing system. This has led to challenges to phonics-centric instruction, with arguments that English orthography encodes morphological and etymological information often obscured by phonics alone, advocating instead for teaching the logic of the spelling system (beyond_bowers_2017). However, structured phonics approaches using word boxes and word sort procedures have been proven effective in facilitating both word recognition and spelling development (facilitating_joseph_2002), indicating that systematic instruction remains valuable even as the field recognizes the limitations of purely phonological approaches.
@@
Beyond specific intervention techniques, broader pedagogical frameworks shape how spelling instruction is delivered and how learners engage with orthographic knowledge. Inquiry-based learning positions learners as active constructors of knowledge through systematic investigation rather than passive recipients of instruction. This approach follows five distinct phases that characterize the inquiry cycle: orientation, conceptualization, investigation, conclusion, and discussion (phases_pedaste_2015). Theoretical foundations align with the contemporary understanding of active knowledge construction (inquirybased_friesen_2013), although empirical evidence reveals essential nuances. Meta-analytical findings show moderate positive effects on learning outcomes, with instructional guidance playing a critical role in determining effectiveness (metaanalysis_lazonder_2016). Unguided discovery learning proved to be less effective than structured inquiry with appropriate scaffolding, indicating that learner autonomy must be balanced with systematic support. Application to language pedagogy shows that inquiry-based teaching develops a deeper linguistic understanding compared to traditional instruction (inquirybased_lee_2014), but substantial barriers to implementation persist. Institutional constraints, time limitations, and teachers’ resistance to constructivist approaches create friction in classroom contexts (challenges_dorier_2013). These challenges are amplified for learners with spelling difficulties who need additional scaffolding to balance open-ended exploration with systematic investigation of orthographic principles.
@@ 2.1. Intervention Methods
Intervention approaches for struggling spellers have predominantly focused on phonological processing and phonics-based instruction, building on the established relationship between phonological awareness and reading acquisition. Intensive remedial instruction targeting phonological processing can prevent reading failure in young children (preventing_torgesen_1999; intensive_torgesen_2001), with explicit systematic phonics instruction producing better outcomes for at-risk children compared to implicit methods (role_foorman_1998). Early literacy support programs that combine phonological training with reading practice have shown effectiveness (evidence_hatcher_2006), and systematic reviews confirm that phonics training benefits English-speaking poor readers, although the effect sizes remain modest (phonics_mcarthur_2018). Individual response patterns vary considerably, and long-term outcomes depend heavily on the intensity and duration of the intervention. Multisensory instructional approaches have shown particular promise, with significant improvements documented when visual, auditory, and kinesthetic modalities are integrated (impact_schlesinger_2017). Color-coded onset-rime decoding interventions have proven effective for first-grade students at serious risk for reading disabilities (effectiveness_hines_2009). However, critical reviews note that despite advances in understanding dyslexia, the effectiveness of treatment remains variable, with many struggling readers requiring sustained and intensive intervention (current_alexander_2004). Evidence-based interventions must create a "virtuous circle" where gains in one area support the development of related skills (evidencebased_snowling_2011).
@@
Structured Word Inquiry offers a fundamentally different instructional approach that recognizes English spelling as a morphophonemic system where spellings represent sound, meaning, and history in an orderly manner. Rather than treating English orthography as a collection of exceptions and irregular rules, SWI demonstrates that the spelling system becomes highly consistent when words are contextualized within their morphology, phonology, etymology, and definition. Bowers and Kirby (effects_bowers_2010) found that morphological instruction produces significant effects on vocabulary acquisition, establishing that teaching morphemic structure supports both spelling and comprehension. This approach engages children in generating and testing hypotheses about how the spelling system works, helping them understand why words are spelled the way they are, an investigative process that aligns with inquiry-based learning principles discussed previously. Colenbrander et al. (assessing_colenbrander_2021) conducted a randomized controlled trial assessing the effectiveness of SWI for students in grades 3 and 5 with reading and spelling difficulties, demonstrating improved performance on both trained and untrained words.
@@
In practice, SWI instruction is typically delivered by speech-language pathologists working with small groups of students, providing the individualized attention and expert guidance necessary for effective implementation. However, the growing demand for specialized literacy support has created significant shortages of qualified SLPs (addressing_squires_2013), limiting access to these evidence-based interventions. Colenbrander et al.’s (assessing_colenbrander_2021) findings revealed that SWI instruction delivered by teaching assistants with limited training produces significantly less effective outcomes compared to expert-delivered instruction, underscoring the expertise gap and the need for alternative delivery methods that maintain instructional quality while addressing scalability challenges.
@@ 2.2. Digital Tools for Teaching Language
Technological interventions for spelling and vocabulary instruction have evolved along several trajectories, each addressing different aspects of literacy support for children with dyslexia. One line of work has explored tangible interfaces to reduce cognitive load through concrete representations of abstract linguistic concepts. These systems demonstrated that physical manipulatives support young language learners through embodied learning experiences (design_fan_2018; spellbound_pandey_2011). This approach has been adapted for Chinese learners (character_fan_2019), addressing the distinct challenges of character-based writing systems where visual complexity and stroke order play central roles.
Adaptive spelling correction systems represent a second trajectory, accommodating characteristic dyslexic error patterns through specialized spellcheckers trained on dyslexic error corpora (spellchecker_rello_2015; computerbased_rello_2015). These systems learn individual error patterns to provide personalized correction suggestions (polispell_li_2013), representing significant advances in accommodating dyslexic spelling patterns. However, they remain fundamentally corrective rather than instructional, helping users spell correctly without developing a deeper orthographic understanding.
Gamification has emerged as a third approach, using game mechanics to maintain engagement during spelling practice (gamification_gooch_2016). Game-based systems have been developed to address spelling difficulties in different linguistic contexts (game_rauschenberger_2015; lexilearn_weerasinghe_2025; gamification_dymora_2019). Comprehensive reviews of these technologies note that most remain limited to phonetic-based approaches (technologies_rauschenberger_2019), failing to capture the comprehensive inquiry-based methodology that characterizes effective SWI instruction. Context-aware writing assistance has shown promise in improving outcomes (design_wu_2019), yet even sophisticated adaptive systems lack frameworks to support hypothesis generation and testing that make SWI effective.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789137isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Interactive systems and toolsccs: Applied computing Educationccs: Computing methodologies Artificial intelligence
Figure 1. This figures illustrates how SPIRE identifies misspellings (left), generates a program for intervention (center) and subsequently renders them in the interface (right) for learner’s inquiry to surface word related insights
@@ 1. Introduction
School-age students often learn spelling by memorization (Hegland, 2021). However, the way in which a word is spelled is more than just a convention of letters. It reflects patterns of meaning, sound, structure, and spelling conventions that connect words to one another and to their histories. For instance, consider the word ‘reach’ meaning to extend or stretch out. A majority of us can spell this word with ease, and are unlikely to spell it as say <reech>, in which the long /i/ sound is mapped to the grapheme <ee>. In fact, English represents the /i/ sound in many ways, including <ee> as in see, <ea> as in reach, <ie> as in piece, and <e-e> as in these. Memorization helps us recall the correct form. However, for many learners, especially children with language-based learning disorders, rote memorization is insufficient because it recruits only one aspect of the lexical quality of the word (phonology or orthography/spelling) (Perfetti, 2007; Seidenberg and McClelland, 1989). What supports their learning is understanding the logic behind these alternatives; for example, noticing that ‘reach’ and ‘teach’ both use the <ea> to represent /iː/, and that this pattern traces back to the Old English reccan (Harper, 2001). Seen this way, spelling becomes less an arbitrary code to memorize and more a system of patterns that can be reasoned about, making the language more learnable and meaningful (Real Spelling., 2025). In fact, learners who struggle with spelling through memorization are often referred to speech and language therapy. In therapy, speech-language pathologists (SLPs) use inquiry as an approach to help learners make sense of how words are spelled and built (Bowers, 2006). For example, if a learner writes <reech> for “reach”, the SLP may begin by asking how the learner decided on that spelling. If the learner explains that /i/ is often spelled <ee> as in “see” or “speech”, the SLP can validate this reasoning, then introduce a contrast, such as “teach” or “teacher”, to highlight the alternative <ea> spelling. With a different error, such as using <sine> for “sign”, the SLP may draw attention to the silent <g> and connect it to morphologically related words like “signal” or “signature”. This practice, known as Structured Word Inquiry (SWI) (Bowers and Kirby, 2010), turns spelling into a process of investigation rather than memorization, drawing on phonology (the relationship between sounds and letters), semantics (meaning of the word), morphology (the structure and meaning of word parts such as roots, prefixes, and suffixes), and etymology (the historical origins and evolution of words) to explain why words are spelled the way they are. Through these inquiry-based interventions, SLPs help learners build durable knowledge that extends beyond a single correction to a broader understanding of how written language works. Consequently, they learn how to independently investigate words.
@@
In contrast, everyday writing tools take a much narrower approach to spelling errors. A misspelling is detected, underlined, and replaced with the most probable correction. Whether a learner types “reech” because they overgeneralized a phonics rule, misapplied a morphological pattern, or simply hit the wrong key, the system’s response is the same. This transactional design supports error-free text, but it bypasses the chance to engage learners in noticing patterns or testing hypotheses. This is a missed opportunity because therapy time is scarce (Lewis et al., 2025) and spelling occurs within these tools in the context of authentic writing tasks, such as science reports and essay writing. Without support in these everyday contexts, the carryover of SWI pedagogy from therapy to classroom writing is limited. What is needed is a way to embed the principles of SWI into the tools learners already use, turning each correction into a brief, situated learning experience. This motivates our central research question: How might we bring the rich, inquiry-based pedagogy of SLPs into everyday writing tools such that spelling errors become an opportunity to learn about the writing system?
@@
Our contributions include: (1) a Pedagogical Program Synthesis that represents SLP’s instructional moves in a domain-specific language and assembles conditional programs from learner errors in real-time, (2) SPIRE, a spell-checker that embeds the synthesized programs into lightweight, interactive interfaces, and (3) an empirical evaluation with experts and learners demonstrating how SPIRE aligns with professional practice and integrates into writing workflows.
@@ 2. Related Work
Our work builds on three interconnected bodies of research: spelling development and learning difficulties in school-age students, evidence-based interventions designed to support struggling spellers, and recent efforts leveraging AI to deliver adaptive literacy instruction.
@@
Vellutino et al. (Vellutino et al., 2004) synthesized four decades of empirical work, establishing that difficulties in phonological awareness, rapid naming, and verbal working memory constitute core deficits that persist in development. Longitudinal studies of upper elementary students with reading or spelling disabilities reveal how these difficulties in phonological information processing evolve throughout formal schooling (Schmidt et al., 2021). However, the presentation of symptoms varies considerably between individuals, requiring adaptive intervention approaches (Kuerten et al., 2019). Corpus-based documentation allowed for an empirical understanding of characteristic dyslexic error types. DysList systematically cataloged error patterns in multiple languages (Rello et al., 2014), while subsequent analysis identified specific phoneme-grapheme correspondence failures that persist despite repeated exposure to high-frequency vocabulary (Cidrim et al., 2024). These difficulties extend beyond traditional literacy contexts into digital environments, where searchers with dyslexia exhibit distinct information-seeking behaviors requiring specialized interface considerations (Morris et al., 2018).
@@
Ehri’s phase model (Ehri, 1995, 2005) describes progression through distinct developmental stages: pre-alphabetic, partial alphabetic, full alphabetic, and consolidated alphabetic, each characterized by qualitatively different approaches to word recognition. This framework establishes grapheme-phoneme knowledge as the basis for reading acquisition in alphabetic languages (Ehri, 1998). However, Silliman et al. (Silliman et al., 2006) demonstrated that preadolescents with atypical language skills show distinct error patterns reflecting deficits in phonological, morphological, and orthographic dimensions, indicating that effective intervention must address multiple representational systems simultaneously.
@@
Müller et al. (Müller et al., 2020) found that syllable-based reading instruction produces significant improvements in both word-reading accuracy and comprehension in second-grade students.
@@
Meta-analytical findings demonstrate more severe reading impairments in deep orthographies such as English compared to shallow orthographies with consistent grapheme-phoneme mappings (Carioti et al., 2021), suggesting that intervention approaches must account for the characteristics of the writing system. This has led to challenges to phonics-centric instruction, with arguments that English orthography encodes morphological and etymological information often obscured by phonics alone, advocating instead for teaching the logic of the spelling system (Bowers and Bowers, 2017). However, structured phonics approaches using word boxes and word sort procedures have been proven effective in facilitating both word recognition and spelling development (Joseph, 2002), indicating that systematic instruction remains valuable even as the field recognizes the limitations of purely phonological approaches.
@@
Beyond specific intervention techniques, broader pedagogical frameworks shape how spelling instruction is delivered and how learners engage with orthographic knowledge. Inquiry-based learning positions learners as active constructors of knowledge through systematic investigation rather than passive recipients of instruction. This approach follows five distinct phases that characterize the inquiry cycle: orientation, conceptualization, investigation, conclusion, and discussion (Pedaste et al., 2015). Theoretical foundations align with the contemporary understanding of active knowledge construction (Friesen and Scott, 2013), although empirical evidence reveals essential nuances. Meta-analytical findings show moderate positive effects on learning outcomes, with instructional guidance playing a critical role in determining effectiveness (Lazonder and Harmsen, 2016). Unguided discovery learning proved to be less effective than structured inquiry with appropriate scaffolding, indicating that learner autonomy must be balanced with systematic support. Application to language pedagogy shows that inquiry-based teaching develops a deeper linguistic understanding compared to traditional instruction (Lee, 2014), but substantial barriers to implementation persist. Institutional constraints, time limitations, and teachers’ resistance to constructivist approaches create friction in classroom contexts (Dorier and García, 2013). These challenges are amplified for learners with spelling difficulties who need additional scaffolding to balance open-ended exploration with systematic investigation of orthographic principles.
@@ 2.1. Intervention Methods
Intervention approaches for struggling spellers have predominantly focused on phonological processing and phonics-based instruction, building on the established relationship between phonological awareness and reading acquisition. Intensive remedial instruction targeting phonological processing can prevent reading failure in young children (Torgesen et al., 1999, 2001), with explicit systematic phonics instruction producing better outcomes for at-risk children compared to implicit methods (Foorman et al., 1998). Early literacy support programs that combine phonological training with reading practice have shown effectiveness (Hatcher et al., 2006), and systematic reviews confirm that phonics training benefits English-speaking poor readers, although the effect sizes remain modest (McArthur et al., 2018). Individual response patterns vary considerably, and long-term outcomes depend heavily on the intensity and duration of the intervention. Multisensory instructional approaches have shown particular promise, with significant improvements documented when visual, auditory, and kinesthetic modalities are integrated (Schlesinger and Gray, 2017). Color-coded onset-rime decoding interventions have proven effective for first-grade students at serious risk for reading disabilities (Hines, 2009). However, critical reviews note that despite advances in understanding dyslexia, the effectiveness of treatment remains variable, with many struggling readers requiring sustained and intensive intervention (Alexander and Slinger-Constant, 2004). Evidence-based interventions must create a "virtuous circle" where gains in one area support the development of related skills (Snowling and Hulme, 2011).
@@
Structured Word Inquiry offers a fundamentally different instructional approach that recognizes English spelling as a morphophonemic system where spellings represent sound, meaning, and history in an orderly manner. Rather than treating English orthography as a collection of exceptions and irregular rules, SWI demonstrates that the spelling system becomes highly consistent when words are contextualized within their morphology, phonology, etymology, and definition. Bowers and Kirby (Bowers and Kirby, 2010) found that morphological instruction produces significant effects on vocabulary acquisition, establishing that teaching morphemic structure supports both spelling and comprehension. This approach engages children in generating and testing hypotheses about how the spelling system works, helping them understand why words are spelled the way they are, an investigative process that aligns with inquiry-based learning principles discussed previously. Colenbrander et al. (Colenbrander et al., 2021) conducted a randomized controlled trial assessing the effectiveness of SWI for students in grades 3 and 5 with reading and spelling difficulties, demonstrating improved performance on both trained and untrained words.
@@
In practice, SWI instruction is typically delivered by speech-language pathologists working with small groups of students, providing the individualized attention and expert guidance necessary for effective implementation. However, the growing demand for specialized literacy support has created significant shortages of qualified SLPs (Squires, 2013), limiting access to these evidence-based interventions. Colenbrander et al.’s (Colenbrander et al., 2021) findings revealed that SWI instruction delivered by teaching assistants with limited training produces significantly less effective outcomes compared to expert-delivered instruction, highlighting the importance of pedagogical fidelity in any delivery method. This expertise gap motivates investigation of whether computational systems can faithfully operationalize the reasoning processes that make expert SWI instruction effective, a prerequisite for eventual scalability that we address through expert evaluation of pedagogical validity.
@@ 2.2. Digital Tools for Teaching Language
Technological interventions for spelling and vocabulary instruction have evolved along several trajectories, each addressing different aspects of literacy support for children with dyslexia. One line of work has explored tangible interfaces to reduce cognitive load through concrete representations of abstract linguistic concepts. These systems demonstrated that physical manipulatives support young language learners through embodied learning experiences (Fan et al., 2018; Pandey and Srivastava, 2011). This approach has been adapted for Chinese learners (Fan et al., 2019), addressing the distinct challenges of character-based writing systems where visual complexity and stroke order play central roles.
Adaptive spelling correction systems represent a second trajectory, accommodating characteristic dyslexic error patterns through specialized spellcheckers trained on dyslexic error corpora (Rello et al., 2015a, b). These systems learn individual error patterns to provide personalized correction suggestions (Li et al., 2013), representing significant advances in accommodating dyslexic spelling patterns. However, they remain fundamentally corrective rather than instructional, helping users spell correctly without developing a deeper orthographic understanding.
Gamification has emerged as a third approach, using game mechanics to maintain engagement during spelling practice (Gooch et al., 2016). Game-based systems have been developed to address spelling difficulties in different linguistic contexts (Rauschenberger et al., 2015; Weerasinghe et al., 2025; Dymora and Niemiec, 2019). Comprehensive reviews of these technologies note that most remain limited to phonetic-based approaches (Rauschenberger et al., 2019), failing to capture the comprehensive inquiry-based methodology that characterizes effective SWI instruction. Context-aware writing assistance has shown promise in improving outcomes (Wu et al., 2019), yet even sophisticated adaptive systems lack frameworks to support hypothesis generation and testing that make SWI effective.